# Linguistic Collapse: Neural Collapse in (Large) Language Models

### v1

@@ Abstract
Neural collapse ($\mathcal{NC}$) is a phenomenon observed in classification tasks where top-layer representations collapse into their class means, which become equinorm, equiangular and aligned with the classifiers. These behaviors – associated with generalization and robustness – would manifest under specific conditions: models are trained towards zero loss, with noise-free labels belonging to balanced classes, which do not outnumber the model’s hidden dimension. Recent studies have explored $\mathcal{NC}$ in the absence of one or more of these conditions to extend and capitalize on the associated benefits of ideal geometries. Language modeling presents a curious frontier, as training by token prediction constitutes a classification task where none of the conditions exist: the vocabulary is imbalanced and exceeds the embedding dimension; different tokens might correspond to similar contextual embeddings; and large language models (LLMs) in particular are typically only trained for a few epochs. This paper empirically investigates the impact of scaling the architectures and training of causal language models (CLMs) on their progression towards $\mathcal{NC}$. We find that $\mathcal{NC}$ properties that develop with scaling are linked to generalization. Moreover, there is evidence of some relationship between $\mathcal{NC}$ and generalization independent of scale. Our work therefore underscores the generality of $\mathcal{NC}$ as it extends to the novel and more challenging setting of language modeling. Downstream, we seek to inspire further research on the phenomenon to deepen our understanding of LLMs – and neural networks at large – and improve existing architectures based on $\mathcal{NC}$-related properties.
@@ 1.1 Neural Collapse
A learning phenomenon known as neural collapse ($\mathcal{NC}$) emerges during the terminal phase of training neural networks with cross-entropy (CE) loss for classification [1]. It was originally characterized as the co-occurrence of the following properties in a model’s top-layer features (also known as last-layer representations or embeddings) and classifiers:
@@
    Within-class variability collapse: Top-layer features collapse to their class means.
@@
    Convergence to self-duality: Top-layer classifier vectors converge to their corresponding class means, and thus also form a simplex ETF.
@@
    Nearest decision rule: Linear classifiers approximate a nearest-class center (NCC) classifier: top-layer embeddings predict the class with the closest mean (implied by $\mathcal{NC}$1-3).
These behaviors, often associated with improved generalization and robustness (among other benefits, such as those discussed in §1.4), traditionally manifest under the following conditions:
  1. C1)
@@ 3. C3)
  4. C4)
    Terminal phase of training (TPT): The model is trained past zero error, towards zero loss.
@@ 4. C4)
To address these, we first examine the specific settings of training CLMs as they are opposed ($\neg$) to the prerequisites (C1-4, §1.1) for $\mathcal{NC}$ to manifest.
  1. $\neg$¬C1)
@@ 1.4 Significance
Recently, methods building on $\mathcal{NC}$ have found use in deep learning at large. We highlight areas such as federated learning [17], graph neural networks [18], incremental/continual learning [19, 20, 21], meta-learning [21, 22], out-of-distribution detection [23, 24, 25], privacy [26, 27], learning theory [28, 29, 30, 31, 32, 33] transfer learning [34, 35, 36, 37, 38, 39], and even LLM prompting [40]. With our results, we aim to extend insights from such contributions and other related works to the language modeling domain and ultimately assist researchers in improving and interpreting their (L)LMs.
@@ 2 Related Works
$\mathcal{NC}$ was initially observed in image classification tasks such as on CIFAR-10/100 [41] and ImageNet [42]. Since then, the phenomenon has been further studied both theoretically and empirically [15, 32, 37, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65], with several works venturing into settings without some of the traditional prerequisites (C1-4, §1.1):
A large number of classes
$\mathcal{NC}$ established the simplex ETF as an optimal configuration. However, a perfect simplex ETF ($\mathcal{NC}_{2}$) requires that the number of classes $C$ not exceed $d+1$ where $d$ is the embedding dimension. This requirement that $d$ be sufficiently large is not practical when the classes number beyond the thousands. GPT-2 [66], for instance, has a vocabulary of over 50K tokens.  [^2] Following [67], one might describe such a setting ($C>d+1$) as a model “in superposition”.
In such a scenario, one might still expect class means to be uniformly distributed on a $d$-dimensional hypersphere [15]. [16] formalize this as hyperspherical uniformity within a generalized neural collapse ($\mathcal{GNC}$) framework, which [6] then empirically demonstrate. These latter two works mention language modeling as applicable for $\mathcal{GNC}$; [6] even frame token prediction as a classification task, just as we do. We remark however that both earlier works simulate a large number of classes by drastically shrinking the embedding dimension. In contrast, we study next-token prediction, using the full class space (vocabulary) with an imbalanced token distribution and ambiguous samples.
Class imbalance
@@
Multi-label classification
In some problems, one might encounter mixed or multi-label samples, be they natural or due to noise or augmentation [69, 70]. $\mathcal{NC}$ was also recently studied for such data by [63], who observed that multi-label class means arrive at the average of their individual labels’ means. They also devise an augmented CE loss function to accommodate such samples.
However, most of our ambiguous samples are not soft- nor multi-label, but rather identical (or very similar) context samples with different hard token labels ($\neg$¬C3, §1.2). This effectively rules out the prospect of achieving zero classification error and potentially introduces irreducible noise.
@@
Early stages of training
[52] studied $\mathcal{NC}$ in small ResNet [71] models that had not yet converged (similar to most LLMs). They show that within-class variability drops and “plateaus” ($\mathcal{NC}$1) earlier than other $\mathcal{NC}$ metrics, a result that we also observe (§4.1, Figures 5, 14).
@@
A more related work studied feature collapse in individual word representations [76], but the authors note that their analysis is limited to shallow NLP modeling on more rigid and symmetrical text data.
@@ 3 Methodology
In this section, we describe the training setup for our CLMs (§3.1), procedures for collecting top-layer embeddings (§3.2, 3.7), and measurements of $\mathcal{NC}$ and generalization (§3.4, 3.5, 3.6, 3.7, 3.8).
@@ 3.1 Dataset and Language Models
TinyStories is a synthetic333 TinyStories was developed and evaluated as a faithful emulation of large language modeling [2], so we chose it for experimentation to train hundreds of CLMs and analyze embeddings at low cost. See Appendix B. dataset generated by GPT-3.5 and GPT-4 [2] with token vocabulary $\mathbb{V}=\llbracket 1,29233\rrbracket$. Token prediction can therefore be framed as classification among $C=$29,233$$ token classes. [^4] Although the GPT-2 [77] tokenizer has over 50K tokens, only a subset vocabulary appears in TinyStories. Following the GPT-style regime[5], raw sequences are packed into $S$ chunks of size $T$, providing $N=S(T-1)$ token samples for training. [^5] We cannot use the first ground-truth nor last predicted token in any chunk. Details are listed in Appendix B.
We use 30 CLM architectures based on GPT Neo [78], configured similarly to [2]. They vary in width (embedding dimension) $d\in\{64,128,256,512,768,1024\}$ and depth (number of self-attention layers) $L\in\{1,2,4,8,12\}$. Our models were trained by teacher-forcing [^6] Parallelized training using sequences’ ground-truth labels for context as opposed to predicted tokens. using CE loss. For each architecture, we trained multiple models for 1, 3, and 10 epochs. We albated over weight decay factors $\beta=0.0005$ [50] and $\beta=0.1$ [77]. Further details are listed in Appendices A, C.
@@ 3.2 Context Embeddings
Base CLMs perform next-token prediction: given a sequence of tokens $\boldsymbol{x}_{1:t}\in\mathbb{V}^{t}$, a top-layer context embedding $\boldsymbol{h}(\boldsymbol{x}_{1:t})\in\mathbb{R}^{d}$ is used to predict the next token $x_{t+1}^{\prime}\in\mathbb{V}$ where $1\leq t\leq T$. A classifier for class $c$ with weights $\boldsymbol{w}_{c}$ and bias [^7] Similar to many CLMs [78, 79, 80, 81], our classifiers do not include additive biases, so $b_{c}=0$. $b_{c}$ would make maximum a prior (MAP) predictions as
@@
For each token class $c$, we are interested in the mean embedding $\boldsymbol{\mu}_{c}\in\mathbb{R}^{d}$ across sequences $s$ and their contexts $\boldsymbol{x}_{1:t}^{(s)}$, where the next token $x_{t+1}^{(s)}$ is ground-truth ($t<T$) and equal to $c$. To center these means, we compute their unweighted  [^8] Different from most literature where classes were balanced and weighting is already equal. average $\boldsymbol{\bar{\mu}}\in\mathbb{R}^{d}$. Precisely, we take
@@
In a second pass, we accumulate the unbiased sample variances [^9] This sample variance is computed across all unnormalized dimensional entries.:
$\sigma_{c}^{2}:=\frac{1}{N_{c}-1}\sum_{s=1}^{S}\sum_{t=1}^{T-1}\left\|\boldsymbol{h}\left(\boldsymbol{x}_{1:t}^{(s)}\right)-\boldsymbol{\mu}_{c}\right\|_{2}^{2}\mathbb{I}\left(x_{t+1}^{(s)}=c\right).$ (3)
@@ 3.3 Homogeneity and Variability
For some collapse measures (such as $\mathcal{(G)NC}_{2}$ and $\mathcal{NC}_{3}$), we are primarily interested in the variation rather than the average of pair-wise relationships. To that end, we also include in our analysis the coefficient of variation (CoV) of several measurements, which is the ratio of their standard deviations to their means: $\sigma(\cdot)/\mu(\cdot)$. We can interpret this as a normalized measure of variability.
3.4 Signal-to-Noise Ratio – $\mathcal{NC}_{1}$
The ability to disambiguate between classes depends on the ratio of within-class to between-class variabilities. Building upon foundational works [82, 83], $\mathcal{NC}$ originally measured variability through an inverse signal-to-noise ratio (SNR), whose minimization constitutes within-class variability collapse ($\mathcal{NC}_{1}$). We employ the class-distance normalized variance (CDNV) proposed by [34]:
$\text{CDNV}_{c,c^{\prime}}:=\frac{\sigma_{c}^{2}+\sigma_{c^{\prime}}^{2}}{2\|\boldsymbol{\mu}_{c}-\boldsymbol{\mu}_{c^{\prime}}\|_{2}^{2}},\quad\forall c\neq c^{\prime}.$ (4)
These pair-wise measures constitute the off-diagonal [^10] The main diagonal of CDNVs is undefined (due to the zero denominator) and ignored. entries of a symmetric matrix in $\mathbb{R}^{C\times C}$, whose average we use as an inverse SNR. Within-class variability collapse is then re-characterized by the minimization of this quantity: $\text{CDNV}_{c,c^{\prime}}\rightarrow 0,\forall c\neq c^{\prime}$. This alternative convergence is faithful to $\mathcal{NC}_{1}$ but more robust and numerically stable [34].
3.5 Geometric Structures – $\mathcal{(G)NC}_{2}$
@@
$\mathcal{NC}_{2}$ also entails that class means are equiangular about their center $\boldsymbol{\bar{\mu}}$: pair-wise distances and angles between their class means should be maximized and similar. Following [1], we measure interference (sometimes known as similarity or coherence [86, 87]). Its minimization,
@@
together with equinormness (Equation 5) constitute convergence to a simplex ETF. Although this geometry is not ultimately attainable since there are too many classes ($C>d+1$), it can still be meaningful to measure a model’s tendency towards one. As with CDNV noise (Equation 4), pair-wise interferences form off-diagonal [^11] The main diagonal of interferences is equal to $\boldsymbol{1}$ (maximal coherence or self-similarity). entries in a symmetric matrix in $\mathbb{R}^{C\times C}$. The minimization of CoV in interferences therefore expresses the degree of “equiangularity”.
Hypershperical uniformity
A relaxation from the ETF is hyperspherical uniformity ($\mathcal{GNC}_{2}$), with mean vectors $\boldsymbol{\mu}_{c}$ uniformly distributed on the $d$-dimensional hypersphere [15, 16]. We gauge the degree of this uniformity with pair-wise interactions under a logarithmic inverse distance kernel:  [^12] Following [16], we employ the logarithmic inverse distance kernel $K_{\log}(\boldsymbol{a},\boldsymbol{b})=\log\|\boldsymbol{a}-\boldsymbol{b}\|^{-1}$ for its ability to emphasize gaps between small distances while scaling down the effect of larger distances.
@@
3.6 Alignment Between Classifiers and Class Mean Embeddings – $\mathcal{(U)NC}_{3}$
The linear classifiers $\{\boldsymbol{w}_{c}\}_{c=1}^{C}$ lie in a dual vector space to that of the class means $\{\boldsymbol{\mu}_{c}\}_{c=1}^{C}$. While convergence to self-duality ($\mathcal{NC}_{3}$) was previously measured as distances [1] between class means classifiers (Equation 11), we instead simplify ($\mathcal{NC}_{3}$) as class-wise similarities:

### v2

@@ Abstract
Neural collapse ($\mathcal{NC}$) is a phenomenon observed in classification tasks where top-layer representations collapse into their class means, which become equinorm, equiangular and aligned with the classifiers. These behaviors – associated with generalization and robustness – would manifest under specific conditions: models are trained towards zero loss, with noise-free labels belonging to balanced classes, which do not outnumber the model’s hidden dimension. Recent studies have explored $\mathcal{NC}$ in the absence of one or more of these conditions to extend and capitalize on the associated benefits of ideal geometries. Language modeling presents a curious frontier, as training by token prediction constitutes a classification task where none of the conditions exist: the vocabulary is imbalanced and exceeds the embedding dimension; different tokens might correspond to similar contextual embeddings; and large language models (LLMs) in particular are typically only trained for a few epochs. This paper empirically investigates the impact of scaling the architectures and training of causal language models (CLMs) on their progression towards $\mathcal{NC}$. We find that $\mathcal{NC}$ properties that develop with scale (and regularization) are linked to generalization. Moreover, there is evidence of some relationship between $\mathcal{NC}$ and generalization independent of scale. Our work thereby underscores the generality of $\mathcal{NC}$ as it extends to the novel and more challenging setting of language modeling. Downstream, we seek to inspire further research on the phenomenon to deepen our understanding of LLMs – and neural networks at large – and improve existing architectures based on $\mathcal{NC}$-related properties. Our code is hosted on GitHub: https://github.com/rhubarbwu/linguistic-collapse.
@@ 1.1 Neural Collapse
A learning phenomenon known as neural collapse ($\mathcal{NC}$) emerges during the terminal phase of training (TPT) deep neural networks with cross-entropy (CE) loss for classification tasks.[1] It was originally characterized as the co-occurrence of the following properties in a model’s top-layer (also known as last-layer) representations (also known as features or embeddings) and linear classifier weights:
@@
    Within-class variability collapse: Top-layer representations collapse to their class means.
@@
    Convergence to self-duality: Linear classifier weight vectors converge to their corresponding top-layer class mean vectors, and thus also form a simplex ETF.
@@
    Nearest decision rule: Linear classifiers approximate a nearest-class center (NCC) classifier: top-layer representations predict the class with the closest mean (implied by $\mathcal{NC}$1-3).
These behaviors, often associated with improved generalization and robustness [3, 4, 5] (among other benefits, such as those discussed in §1.4), traditionally manifest under the following conditions:
  1. Rq1)
@@ 3. Rq3)
  4. Rq4)
    Sufficient training (TPT): The model is trained past zero error, towards zero loss.
@@ 4. Rq4)
To address these, we first examine the specific settings of training CLMs as they are opposed ($\neg$) to the traditional prerequisites (R1-4, §1.1) for $\mathcal{NC}$ to manifest.
  1. $\neg$¬Rq1)
@@ 1.4 Significance
Recently, methods building on $\mathcal{NC}$ have found use in deep learning at large. We highlight areas such as federated learning [20], graph neural networks [21], incremental/continual learning [22, 23, 24], meta-learning [24, 25], out-of-distribution detection [26, 27, 28], privacy [29, 30], learning theory [31, 32, 33, 34, 35, 36] transfer learning [3, 5, 37, 38, 39, 40], and even LLM prompting [41]. With our results, we aim to extend insights from such contributions and other related works to the autoregressive language modeling domain and ultimately assist researchers in improving and interpreting their (L)LMs.
@@ 2 Related Works
$\mathcal{NC}$ was initially observed in image classification tasks such as on CIFAR-10/100 [42] and ImageNet [43]. Since then, the phenomenon has been further studied both theoretically and empirically [5, 18, 35, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66], with several works venturing into settings without some of the traditional prerequisites ($\neg$¬Rq1-4, §1.2) and proposing adaptations of the analysis framework or optimization procedures:
A large number of classes ($\neg$¬Rq1)
$\mathcal{NC}$ established the simplex ETF as an optimal configuration. However, a perfect simplex ETF ($\mathcal{NC}_{2}$) requires that the number of classes $C$ not exceed $d+1$ where $d$ is the embedding dimension. This requirement that $d$ be sufficiently large is not practical when the classes number beyond the thousands. [^2] Following [67], one might describe such a setting ($C>d+1$) as a model “in superposition”. For instance, GPT-2 [68] and Llama 3.1 [69] have vocabularies of around 50K and 128K tokens, respectively.
In such a scenario, one might still expect class means to be uniformly distributed on a $d$-dimensional hypersphere [18]. [19] formalize this as hyperspherical uniformity within a generalized neural collapse ($\mathcal{GNC}$) framework, which [9] then empirically demonstrate. These latter two works mention language modeling as applicable for $\mathcal{GNC}$; [9] even framed token prediction as a classification task, just as we do. We remark however that both earlier works simulate a large number of classes by drastically shrinking the embedding dimension. In contrast, we study next-token prediction, using the full class space (vocabulary) with an imbalanced token distribution and ambiguous samples.
Class imbalance ($\neg$¬Rq2)
@@
Multi-label classification ($\neg$¬Rq3)
In some problems, one might encounter mixed or multi-label samples, be they natural or due to noise or augmentation [71, 72]. $\mathcal{NC}$ was also recently studied for such data by [73], who observed that multi-label class means arrive at the average of their labels’ respective means. They also devise an augmented CE loss function to accommodate such samples.
Likewise, most of our ambiguous samples could be considered multi- or soft-label: identical (or very similar) context samples with different hard token labels ($\neg$¬Rq3). Under popular CLM pre-training (teacher-forcing with CE loss), this effectively precludes the prospect of achieving zero classification error and potentially introduces irreducible noise.
@@
Early stages of training ($\neg$¬Rq4)
[53] studied $\mathcal{NC}$ in small ResNet [74] models that had not yet converged (similar to most LLMs). They show that within-class variability drops and “plateaus” ($\mathcal{NC}_{1}$) earlier than other $\mathcal{NC}$ metrics, a result that we also observe (§4.1, Figures 6, 7).
@@
A more related work studied feature collapse in individual word representations [79], but the authors note that their analysis is limited to shallow NLP modeling on more rigid and tabular text data.
@@ 3 Methodology
Below we describe the training setup for our CLMs (§3.1), procedures [^3] Leveraging our generic $\mathcal{NC}$ package: https://github.com/rhubarbwu/neural-collapse. for collecting top-layer embeddings (§3.2, 3.7), and measurements of $\mathcal{NC}$ and generalization (§3.4, 3.5, 3.6, 3.7, 3.8).
@@ 3.1 Dataset and Language Models
TinyStories [2] is a synthetic444 TinyStories was developed and evaluated as a faithful emulation of large language modeling, so we chose it for experimentation to train hundreds of CLMs and analyze embeddings at low cost. See Appendix A. dataset generated by GPT-3.5 and GPT-4 using around 1500 English words a child might use. Next-token prediction is performed by sampling from the token vocabulary $\mathbb{V}=\llbracket 1,29233\rrbracket$, which for our purposes can therefore be framed as classification among $C=$29,233$$ classes. [^5] Although the GPT-Neo [80] tokenizer has over 50K tokens, only a subset vocabulary appears in TinyStories. Following the GPT-style preprocessing regime [8], raw sequences are packed into $S$ chunks of size $T$, providing $N=S(T-1)$ token samples for training. [^6] We cannot use the first ground-truth nor last predicted token in any chunk. Details are listed in Appendix A.
We use 30 CLM architectures based on GPT Neo [80], configured similarly to [2]. They vary in width (embedding dimension) $d\in\{64,128,256,512,768,1024\}$ and depth (number of self-attention layers) $L\in\{1,2,4,8,12\}$. Our models were trained by teacher-forcing [^7] Parallelized training using sequences’ ground-truth labels for context as opposed to predicted tokens. using CE loss. For each architecture, we trained multiple models for 1, 3, and 10 epochs ablated over weight decay factors $\beta=0.0005$ [51] and $\beta=0.1$ [81]. Further details are listed in Appendices B, C.
@@ 3.2 Context Embeddings
Base CLMs perform next-token prediction: given a sequence of tokens $\boldsymbol{x}_{1:t}\in\mathbb{V}^{t}$, a top-layer context embedding $\boldsymbol{h}(\boldsymbol{x}_{1:t})\in\mathbb{R}^{d}$ is used to predict the next token $x_{t+1}^{\prime}\in\mathbb{V}$ where $1\leq t\leq T$. A classifier for class $c$ with weights $\boldsymbol{w}_{c}$ and bias [^8] Similar to many causal LLMs [80, 81, 82, 83, 84], our classifiers do not include additive biases, so $b_{c}=0$. $b_{c}$ would make maximum a prior (MAP) predictions as
@@
For each token class $c$, we are interested in the mean embedding $\boldsymbol{\mu}_{c}\in\mathbb{R}^{d}$ across sequences $s$ and their contexts $\boldsymbol{x}_{1:t}^{(s)}$, where the next token $x_{t+1}^{(s)}$ is ground-truth ($t<T$) and equal to $c$. To centre these means, we compute their unweighted [^9] Different from most literature where classes were balanced and weighting is already equal. average $\boldsymbol{\bar{\mu}}\in\mathbb{R}^{d}$. Precisely, we take
@@
In a second pass, we accumulate the sample variance norms: [^10] This sample variance is computed across all unnormalized dimensional entries.
$\sigma_{c}^{2}:=\frac{1}{N_{c}}\sum_{s=1}^{S}\sum_{t=1}^{T-1}\left\|\boldsymbol{h}\left(\boldsymbol{x}_{1:t}^{(s)}\right)-\boldsymbol{\mu}_{c}\right\|_{2}^{2}\mathbb{I}\left(x_{t+1}^{(s)}=c\right).$ (3)
@@ 3.3 Homogeneity and Variability
For some collapse measures (such as $\mathcal{(G)NC}_{2}$ and $\mathcal{NC}_{3}$), we are primarily interested in the variation rather than the average of pairwise relationships. To that end, we also include in our analysis the coefficient of variation (CoV) of several measurements, which is the ratio of their standard deviations to their means: $\sigma(\cdot)/\mu(\cdot)$. We can interpret this as a normalized measure of variability.
3.4 Signal-to-Noise Ratio – $\mathcal{NC}1$
The ability to disambiguate between classes depends on the ratio of within-class to between-class variabilities. Building upon foundational works [85, 86], $\mathcal{NC}$ originally measured variability through an inverse signal-to-noise ratio (SNR), whose minimization constitutes within-class variability collapse ($\mathcal{NC}_{1}$). We instead employ a class-distance normalized variance (CDNV) similar to [3]:
${\hat{\sigma}}_{c,c^{\prime}}:=\frac{1}{\|\boldsymbol{\mu}_{c}-\boldsymbol{\mu}_{c^{\prime}}\|_{2}^{k}}\cdot\frac{\sigma_{c}^{2}+\sigma_{c^{\prime}}^{2}}{2\|\boldsymbol{\mu}_{c}-\boldsymbol{\mu}_{c^{\prime}}\|_{2}^{2}},\quad\forall c\neq c^{\prime}.$ (4)
Our metric differs in that we divide by an additional power $\|\boldsymbol{\mu}_{c}-\boldsymbol{\mu}_{c^{\prime}}\|_{2}^{k}$ of the mean distance norm. This further downweights the CDNV within well-separated class pairs in favour of emphasizing more mutually noisy pairs. We found this augmented CDNV with $k=2$ to be especially useful in our setting of many imbalanced and variably confusable token classes.
These pairwise measures constitute the off-diagonal [^11] The main diagonal of CDNVs is undefined (due to the zero denominator) and ignored. entries of a symmetric matrix in $\mathbb{R}^{C\times C}$, whose average we use as an inverse SNR. Within-class variability collapse is then re-characterized by the minimization of this quantity: ${\hat{\sigma}}_{c,c^{\prime}}\rightarrow 0,\forall c\neq c^{\prime}$. This alternative convergence is empirically faithful to $\mathcal{NC}_{1}$ but more robust and numerically stable [3].
3.5 Geometric Structures – $\mathcal{(G)NC}2$
@@
$\mathcal{NC}_{2}$ also entails that class means are equiangular about their center $\boldsymbol{\bar{\mu}}$: pairwise distances and angles between their class means should be maximized and similar. Following [1], we measure interference (sometimes known as similarity or coherence [89, 90]). Its minimization,
@@
together with equinormness (Equation 5) constitute convergence to a simplex ETF. Although this geometry is not ultimately attainable since there are too many classes ($C>d+1$), it can still be meaningful to measure a model’s tendency towards one. As with CDNV noise (Equation 4), pairwise interferences form off-diagonal [^12] The main diagonal of interferences is equal to $\boldsymbol{1}$ (maximal coherence or self-similarity). entries in a symmetric matrix in $\mathbb{R}^{C\times C}$. The minimization of CoV in interferences therefore expresses the degree of “equiangularity”.
Hyperspherical uniformity
A relaxation from the ETF is hyperspherical uniformity ($\mathcal{GNC}_{2}$), with mean vectors $\boldsymbol{\mu}_{c}$ uniformly distributed on the $d$-dimensional hypersphere [18, 19]. We gauge the degree of this uniformity with pairwise interactions under a logarithmic inverse distance kernel: [^13] Following [19], we employ the logarithmic inverse distance kernel $K_{\log}(\boldsymbol{a},\boldsymbol{b})=\log\|\boldsymbol{a}-\boldsymbol{b}\|_{2}^{-1}$ for its ability to emphasize gaps between small distances while scaling down the effect of larger distances.
@@
3.6 Alignment Between Classifiers and Class Mean Embeddings – $\mathcal{(U)NC}3$
The linear classifiers $\{\boldsymbol{w}_{c}\}_{c=1}^{C}$ lie in a dual vector space to that of the class means $\{\boldsymbol{\mu}_{c}\}_{c=1}^{C}$. While convergence to self-duality ($\mathcal{NC}_{3}$) was initially measured as distances [1] between class means and classifiers (Equation 11), we follow [5] to inspect class-wise cosine similarities: [^14] Dot-product would be confounded by norms and therefore inappropriate for similarity up to rescaling.